OpenStudy (anonymous):

Just need a little help understanding how to do these simple set theory proofs

OpenStudy (anonymous):

Questions are here

OpenStudy (anonymous):

So i'm thinking we can use x as an element of A and show that it is not equal for a counter-example, but I haven't done a proof for these before

ganeshie8 (ganeshie8):

\(\overline{A-B} = \overline{A}-\overline{B}\) Counterexample : U = {1, 2, 3} A = {1, 2} B = {2, 3} \(\overline{A-B}=?\) \(\overline{A}-\overline{B}=?\)

OpenStudy (anonymous):

So one would be: A' - B' = {3}?

OpenStudy (anonymous):

and (A - B)' = {2,3}

OpenStudy (anonymous):

I'm not sure about the first one

ganeshie8 (ganeshie8):

{3} is not the same set as {2, 3} so first statement is false

OpenStudy (anonymous):

oh yeah, I see that

OpenStudy (anonymous):

so when proving we can use the same method?

ganeshie8 (ganeshie8):

Counterexample is for disproving a statement

OpenStudy (anonymous):

like in the next one A-B = {1} and B'-A' = {1} is not a valid proof?

ganeshie8 (ganeshie8):

We need to use other techniques for proving

ganeshie8 (ganeshie8):

Showing one example for which the statement is true is NOT a proof.

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Has to hold for all cases right?

ganeshie8 (ganeshie8):

But, Showing one COUNTER example for which the statement is NOT true is sufficient for disproving.

ganeshie8 (ganeshie8):

So, we are good with part a proof. We have proved that the given statement does not hold by providing a counterexample

ganeshie8 (ganeshie8):

Any ideas on part b ?

OpenStudy (anonymous):

can we use laws of logic? like go LHS = RHS?

ganeshie8 (ganeshie8):

Yes, recall below relationship : \[A-B = A\cap \overline{B}\]

OpenStudy (anonymous):

okay so we have to make that look like B' ^ A?

OpenStudy (anonymous):

Oh sorry, you said that

ganeshie8 (ganeshie8):

\[A-B = A\cap \overline{B}\tag{1}\] \[\overline{B}-\overline{A} = \overline{B} \cap \overline{\overline{A} } = \overline{B} \cap A\tag{2} \]

ganeshie8 (ganeshie8):

from (1) and (2), can we conclude \(A-B = \overline{B}-\overline{A}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okie got it :)

OpenStudy (anonymous):

For the second group can we use the same type of counter example?

ganeshie8 (ganeshie8):

For the second group, part a looks like a correct statement

OpenStudy (anonymous):

So logic laws again?

OpenStudy (anonymous):

Kind of looks like a distributive law

ganeshie8 (ganeshie8):

yeah lets do part b first because it is easy

OpenStudy (anonymous):

alrighty

ganeshie8 (ganeshie8):

part b : \(C\cup (A-B) =(C\cup A)-(C\cup B)\) Counterexample : U = {1, 2, 3,4} A = {1, 2} B = {1, 3} C = {1, 4} \(C\cup (A-B)=?\) \((C\cup A)-(C\cup B)=?\)

OpenStudy (anonymous):

first one is = {1,2,4)?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

second one is {2}?

ganeshie8 (ganeshie8):

Yes, they are not same; so, we're done !

OpenStudy (anonymous):

Hooray :D

OpenStudy (anonymous):

so the logic law one now

OpenStudy (anonymous):

we use the A - B = A ^ B' thing again right?

OpenStudy (anonymous):

kind of confused here

OpenStudy (anonymous):

Isn't it supposed to be A - B = A union B' ?

OpenStudy (anonymous):

not intersection?

ganeshie8 (ganeshie8):

Yeah try this : \[C\cap (A-B) = C\cap (A \cap \overline{B})\tag{1}\] \[(C\cap A) - (C\cap B) \\= (C\cap A)\cap \overline{ (C\cap B)} \\=(C\cap A)\cap (\overline{C}\cup \overline{B}) \\= [(C\cap A)\cap \overline{C}] \cup [(C\cap A)\cap \overline{B}]\\=[\emptyset ]\cup [(C\cap A)\cap \overline{B}]\\=(C\cap A)\cap \overline{B} \]

ganeshie8 (ganeshie8):

what can we conclude from (1) and the last line ?

OpenStudy (anonymous):

RHS = LHS

OpenStudy (anonymous):

so true, thanks ganeshie